OpenStudy (anonymous):

Will give medal! Solve by using either elimination. Please show work. 3x – 9y = 3 6x – 3y = -24

OpenStudy (anonymous):

Anything you do to one side of an equation AND to the other will keep it balanced. So your trying to get only one variable. 1. Try multiplying each term in the top equation by three. 2.subtract the bottom equation from the top 3. now your left with just (something)y = (something) 4. solve this to find out what y is 5. put that value of y back into the original equation and solve for x

OpenStudy (anonymous):

When i multiply the terms on top by 3 i get 9x-27y=9 when i subtract the bottom from the top i'll have 2x-30y=15 did you mean multiply by 2 for the terms on top?

OpenStudy (anonymous):

yes by 2 :)

OpenStudy (anonymous):

or the bottom one by 3 in order to eliminate the y terms if you wanted

OpenStudy (anonymous):

oh ok i gotcha

OpenStudy (anonymous):

ok so i got that y=2, am i right so far?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok for my answer i got x=7 is that right? @JamesWolf

OpenStudy (anonymous):

wait no its x=5

OpenStudy (anonymous):

x = -5, y = 2

OpenStudy (anonymous):

sorry, x = -5, y = -2

OpenStudy (anonymous):

ohhh ok yea i saw what i did wrong, didn't carry the negative. Thanks! Hey can help me with one more using a different method?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

y = 1/2x – 6 2x + 6y = 19 Using substitution method

OpenStudy (anonymous):

@JamesWolf

OpenStudy (anonymous):

so y = 1/2x -6 substitute this value of y into the second equation and solve for x that is, where the y is in the other equation ....6y..... Place in (1/2x -6) so you will have 2x + 6(1/2x - 6) = 19 now you only have one variable and you can sollve for x

OpenStudy (anonymous):

ok thanks!